OpenStudy (anonymous):

16y-41=38. I know how to do it but do I divide my answer??

pooja195 (pooja195):

did u add 41?

pooja195 (pooja195):

i know you will end up with a decimal turn it into a fraction :)

OpenStudy (anonymous):

@pooja195 yes I did

pooja195 (pooja195):

what did you get?

OpenStudy (anonymous):

48

pooja195 (pooja195):

incorrect add 41 to38

OpenStudy (anonymous):

Omg I. So sorry I just realized I mixed up another problem with this one that's why I got 48 because it's 16y-41=7 I'm so sorry

pooja195 (pooja195):

haha your fine :D do you need help?

OpenStudy (anonymous):

Yes do I divide or multiply?

pooja195 (pooja195):

add 41

OpenStudy (anonymous):

I did so I multiply or divide my answer?

pooja195 (pooja195):

then you divide by 16 i beleive it will be a decimal

OpenStudy (anonymous):

Ok I got 0.33

OpenStudy (anonymous):

Or 0.333

pooja195 (pooja195):

Simplifying 16y + -41 = 7 Reorder the terms: -41 + 16y = 7 Solving -41 + 16y = 7 Solving for variable 'y'. Move all terms containing y to the left, all other terms to the right. Add '41' to each side of the equation. -41 + 41 + 16y = 7 + 41 Combine like terms: -41 + 41 = 0 0 + 16y = 7 + 41 16y = 7 + 41 Combine like terms: 7 + 41 = 48 16y = 48 Divide each side by '16'. y = 3 Simplifying y = 3

OpenStudy (anonymous):

Thank u <3

pooja195 (pooja195):

your welcome :)

pooja195 (pooja195):

\(\huge\color{Red}{WELCOME~TO~OPENSTUDY!!!}\)